(9x + 3) + (7x + 1)
9x + 3 + 7x + 1= 180
16x + 4= 180
16x + 4 - 4= 180 - 4
16x= 176
x= 11
I not sure how to work 9x + 3 AND 7x + 1= measure of the larger angle and measure of the smaller angle= too get the answer correct.

Answer by ankor@dixie-net.com(22740)

(9x + 3) + (7x + 1)
9x + 3 + 7x + 1= 180
16x + 4= 180
16x + 4 - 4= 180 - 4
16x= 176
x= 11
:
You've done the main part, just substitute 11 for x in the two original expressions
9(11) + 3 = 102 degrees
7(11) + 1 = 78 degrees
-----------------------
totals to 180 degrees which is what it should do